Find the angle between the vectors -4i+3j+5k and 1i+3j+5k?


A. 48 degree


B. 44.1 degree


C. 24.7 degree


D. 22.1 degree


Grade:Upto college level

3 Answers

BHUKYA ARUN KUMAR
10 years ago

48 degree

mallanna goud
10 years ago

dot product of given vectors=30

magnitudes of vectors -4i+3j+5k and1i+3j+5k are 7.07 , 5.92

7.07*5.92*cosa=30

cosa=30/7.07*5.92

cosa=30/41.85

cosa=0.7167

a=1/cos(0.7167)

a=44.1

treat as *=multiplication and a as angle

sumanth
10 years ago

44.1